# The Open mulTiwavelength Transient Event Repository (OTTER): Infrastructure Release and Tidal Disruption Event Catalog

### v1

@@ 3.1.1 Uploading Data via the OTTER Website
To help maintain (in the case of TDEs) and build (in the case of other transients) a complete dataset in OTTER, we encourage the community to upload their data to OTTER when it is accepted for publication as a standard part of their workflow. The only data we require users to provide is their name, their email, the transient name, the transient coordinates, and the bibcode we should reference for that data. Other optional metadata fields include the redshift, luminosity distance, comoving distance, discovery date, and classification. For each piece of metadata the user is required to provide a citable bibcode.
@@ 3.2 OTTER API
As discussed in §2.3, some photometry stored in OTTER may be different reductions of the same dataset. This especially affects X-ray and UV data and is intentional because it provides users the freedom to choose their preferred reduction. So, as part of the query implemented in the Python API, we provide an algorithm for finding different reductions of the same data and choosing only one of those photometry points. The algorithm finds any photometry points within a specified date tolerance ($\sim$ 1 day) of the observation (or between the minimum and maximum date for binned data) of the same transient from the same telescope and filter but from a different reference. Then, to choose the photometry to be returned to the user, we first check if only one of the reductions is host subtracted and, if so, we default to that value. If neither reduction or more than one reduction are host subtracted, then we take the most recent reduction. Our testing reveals that this algorithm works very well for the current OTTER dataset. If the user disagrees with this algorithm, they are instead able to “plug in” their favorite algorithm to the query (for example, an algorithm that always prefers the reduction from one paper over another).
@@
4 Current State of the OTTER Dataset
Figure 7: Cumulative histogram of the number of TDEs discovered per year, colorized by the wavelength range in which it was discovered. The approximate start date of the major TDE discovery telescopes are labelled on the top. Note the logarithmic scale on the y-axis. Figure 16 is a non-cumulative version of this histogram in Appendix C. A movie of this figure with a linear y-axis is available at: https://github.com/astro-otter/examples/blob/main/tde-discovery-histogram-linear.gif
@@ 4 Current State of the OTTER Dataset
Figure 9: Light curves in $\nu L_{\nu}$ space of all of the photometry in OTTER. The colorbars have different minima and maxima based on the number of observations in that wavelength regime and are normalized by the width of the bin in days. Note the two distinct classes of Radio TDE light curves (jetted and non-jetted radio emission). The UV/Optical/IR and X-ray photometry span $\sim 6$ orders of magnitude in luminosity space, demonstrating a diversity of in UVOIR and X-ray TDE luminosities.
@@
Figure 10: Distribution of photometry points in OTTER as a function of Wavelength/Frequency. We have labelled some typical filters along the top for reference. Most of the photometry is in the UVOIR bands and the X-ray and Radio distributions have about the same number of photometry points.
@@
As of 2025 September 1, OTTER contains data on 232 TDEs, or TDE candidates, spanning 34 years of observations, making it the largest catalog of TDE photometry and metadata to date. Figure 7 summarizes the TDE discovery rate by different X-ray, radio, and optical surveys over this time period. It also shows the projection for the TDE discovery rate over the next 5 years as the Rubin Observatory transient discovery rate increases. This is predicted to increase the TDE discovery rate by two orders of magnitude (van Velzen et al., 2011; Bricman & Gomboc, 2020).
@@
For every transient in OTTER, we have at least one set of coordinates. $219/232\approx 94\%$ of the transients in OTTER have a spectroscopic redshift. The transients missing a redshift are MAXI J1807+132, which is an X-ray selected TDE candidate, and the 12 radio selected TDEs from Dykaar et al. (2024). Figure 8 shows the redshift distribution of known TDEs in OTTER, not corrected for observational bias. The peak appears around $z\sim 0.09$ and 68% of TDEs lie in the range from $z\sim 0.03-0.2$, which is likely the result of an observational bias.
@@
$195/232\approx 84\%$ of the transients in OTTER have a discovery date stored. The missing discovery dates are from TDEs that were not discovered on a specific date but were rather found in archival searches (e.g., Masterson et al., 2024; Somalwar et al., 2023a; Dykaar et al., 2024). We hesitate to set the discovery date of these events to the first archival detection since, generally, archival searches have much larger gaps between observations than dedicated surveys. This means it is difficult to interpret the date of the first detection as the discovery date under the typical definition.
@@
All the photometry in OTTER is shown in Figures 9, 10, and 11 and contains $\gtrsim 80,000$ photometry points. $189/232=81\%$ of the TDEs in OTTER have at least one photometry point at any wavelength. $92/232\approx 40\%$ of the TDEs in OTTER have at least one radio observation, $138/232\approx 59\%$ of TDEs in OTTER have at least one UVOIR observation, and $80/232\approx 34\%$ of TDEs in OTTER have at least one X-ray observation. Some of the missing photometry simply demonstrates inconsistencies in the multiwavelength follow up of TDEs. In more recent years, more TDEs were followed up more consistently across the electromagnetic spectrum (e.g., Figure 11).
@@
Note that in years since $\sim 2022$, our photometric completeness begins to fall off and we have no photometry after $2024$. This is shown in Figure 11 and is expected because OTTER only contains published photometry, so the more recently discovered TDEs simply have fewer published observations.
@@
5 Example Applications of OTTER
The OTTER infrastructure has a wide variety of potential applications across the field of transient astronomy, ranging from a quick view of existing public datasets to detailed quantitative comparison with models and locally-stored datasets. Here, we list a few brief examples of applications of the catalog and infrastructure to the TDE population currently stored in OTTER. We provide jupyter notebooks as part of the documentation demonstrating each of these analyses. The analyses done in this section use the OTTER dataset as of 2025 September 1.
[…]
5.1 Example 1: Quantifying the Delayed X-ray Emission in Optically-Discovered TDEs
@@
Guolo et al. (2024) found, qualitatively, that the X-ray emission of a subset of optically-selected TDEs is delayed compared to their UV/Optical emission. To further explore this, we quantify the time delay for a large sample of TDEs. We select all of the TDE candidates in OTTER that have at least one detection and at least two observations in both the X-rays and UV/Optical wavelengths, with a few exceptions. CSS100217 meets these criteria but we remove it from the sample because it is unlikely a TDE (Drake et al., 2011; Kankare et al., 2017). We also exclude the jetted TDEs Sw J1644+57, Sw J2058+05, and AT 2022cmc since the processes producing the UV/Optical and X-ray emission are unlikely to be the same as the rest of the TDEs in this sample (Zauderer et al., 2011; Levan et al., 2011; Berger et al., 2012; Wiersema et al., 2012; Zauderer et al., 2013; Levan et al., 2016; Yang et al., 2016; Auchettl et al., 2017; Eftekhari et al., 2018; Cendes et al., 2021b; Cenko et al., 2012a; Pasham et al., 2015; Auchettl et al., 2017; Brown et al., 2017). Note, we do not consider infrared photometry for this analysis ($\lambda_{\rm eff}\geq 1~{\rm\mu m}$) because the infrared emission, which is often dominated by dust echoes, appears to be unrelated to the early-time UV/Optical emission (van Velzen et al., 2021a, 2023). These cuts resulted in a sample that is entirely optically selected.
@@
Since our sample only includes optically discovered TDEs that were later followed up with X-ray observations, there may be a systematic bias of $\Delta t_{\rm del}$ to positive values greater than the true population $\Delta t_{\rm del}$ distribution. We account for this bias by introducing upper limits when the maximum was the first observation and lower limits when the maximum was the last observation (e.g., see the arrows in 12a). For subsequent statistical analysis, if $t_{\rm X}$ is an upper limit or $t_{\rm UVO}$ is a lower limit then $\Delta t_{\rm del}$ is treated as a lower limit. In contrast, if $t_{\rm X}$ is a lower limit or $t_{\rm UVO}$ is an upper limit then $\Delta t_{\rm del}$ is treated as an upper limit.
@@
To account for the uncertainties and limits on the delay times in our statistical analysis, we fit a Kaplan-Meier estimator (Kaplan & Meier, 1958) to the $\Delta t_{\rm del}$ distribution [^18] To be more specific, we fit the Kaplan-Meier estimator using the uncertainties and/or limits as interval censored data, essentially treating each of these $\Delta t_{\rm del}$ values as a range (which could have a lower limit of -$\infty$ or an upper limit of $\infty$).. The best-fit Kaplan-Meier estimator has a median $\Delta t_{\rm del}$ of $22$ with 16th and 84th quartiles of 3 and 283 days, respectively. We additionally compute a $95\%$ Greenwood confidence interval (Greenwood, 1926) and find that the true population median of $\Delta t_{\rm del}$ lies in the range 32283-228 days. Overall, based on the quartiles of the $\Delta t_{\rm del}$ distribution, we find a conservative upper limit $\Delta t_{\rm del}\lesssim 300$ days, with one exceptional event with an extremely long delay time (PTF09axc [^19] https://otter.idies.jhu.edu/transient/PTF09axc). This indicates that, for most TDEs, we disfavor theoretical stream-stream models that predict longer circularization times than a few hundred days. However, the data are still consistent with either the theoretical circularization time found by Steinberg & Stone (2024) or the theoretical dissipation times presented in Metzger (2022).
If the $\Delta t_{\rm del}$ distribution is interpreted as the circularization time, the wide spread we find in $\Delta t_{\rm del}$ could be further evidence that the circularization time is physically related to the initial period of the stellar orbit (Bonnerot & Stone, 2021), SMBH properties (e.g., mass and spin, Steinberg & Stone, 2024), or nodal precession of the plane of the stellar stream (Guillochon & Ramirez-Ruiz, 2013; Bonnerot & Stone, 2020). If, instead, the time delay is interpreted as the dissipation time of a reprocessing envelope, the spread in the distribution can be interpreted as a viewing angle effect (Dai et al., 2018; Metzger, 2022). Moving forward, a theoretical prediction of this $\Delta t_{\rm del}$ distribution under both models may be necessary to distinguish between them. We defer further investigation of these potential relationships to future work.
5.2 Example 2: The Dusty Environments of TDE Extreme Coronal Line Emitters
@@
As a sample analysis, we plot the peak radio luminosity (a mix of public and private data; Franz, N et al., in prep) against the WISE colors (which we obtain by using the OTTER API to query WISE) for the ECLE-TDEs in Figure 13 and compare these to the non-ECLE-TDEs population. We find that ECLE-TDEs tend to occur in environments with redder dust than the broader TDE population with a high statistical significance ($p\sim 10^{-6}$). By extension, this may indicate that ECLE-TDEs occur in environments with higher density dust [^21] There is also the possibility that this difference instead indicates a different amount of incident flux onto the dust grains, especially in the case of the archival ECLEs that do not have constraints on the start of the observed emission. But, consider that Hinkle et al. (2024) found that the WISE color is also transient in a subset of our ECLE sample, indicating that the dust heating is transient in these events. We defer further discussion disentangling these scenarios to future work as it will likely require a more detailed multiwavelength investigation.. This is a similar finding to Hinkle et al. (2024) but with a larger sample of events. The radio data in this figure will be discussed in more detail in Franz, N et al., in prep.
5.3 Example 3: Easily Modeling TDE Light-Curves
Another advantage of consolidating the TDE data is that it makes it very easy to apply existing modeling codes, such as MOSFiT (Guillochon et al., 2018; Mockler et al., 2019), redback (Sarin et al., 2024), FitTed (Mummery et al., 2024a), and TDEMass (Ryu et al., 2020a, b). By fitting these models to the data we can derive properties of the SMBH, progenitor star, and cirumnuclear environment. Most notably, these models produce a galaxy-independent measure of the SMBH mass which is either constrained by the light curve as a whole (e.g., MOSFiT) or from the late-time plateau of the light curve (e.g., FitTed).
@@ 5.3 Example 3: Easily Modeling TDE Light-Curves
In this work, we focus on applying MOSFiT to some of the objects in the OTTER dataset. The MOSFiT TDE model begins after the circularization ceases, when accretion onto the SMBH is actively occuring. MOSFiT then assumes that the accretion disk produces thermal X-rays which are reprocessed into the UV/Optical. As an example, we can model two TDEs recently discovered by ZTF: TDE 2020zso (Forster et al., 2020a; Ihanec et al., 2020; Hammerstein et al., 2023; Mummery et al., 2024) and TDE 2021sdu (Munoz-Arancibia et al., 2021a; Chu et al., 2021a; Yao et al., 2023; Mummery et al., 2024). We show these models in Figure 14. From MOSFiT we find that the mass of the SMBH in TDE 2020zso is $\log_{10}\left(M_{BH}\right)=6.14^{+0.13}_{-0.24}$ and of TDE 2021sdu is $\log_{10}\left(M_{BH}\right)=5.53^{+0.22}_{-0.25}$, where the lower and upper uncertainties are the 5% and 95% quantiles, respectively. A jupyter notebook demonstrating how to model these two datasets with MOSFiT is available on the OTTER API documentation webpage and shows that it takes $\lesssim 100$ lines of code to reformat the OTTER data into an input file for MOSFiT.
5.4 Example 4: Validating Optical Transient Classifiers in the Rubin era

### v2

@@ III.1.1 Uploading Data via the OTTER Website
To help maintain (in the case of TDE candidates) and build (in the case of other transients) a complete dataset in OTTER, we encourage the community to upload their data to OTTER when it is accepted for publication as a standard part of their workflow. The only data we require users to provide is their name, their email, the transient name, the transient coordinates, and the bibcode we should reference for that data. Other optional metadata fields include the redshift, luminosity distance, comoving distance, discovery date, and classification. For each piece of metadata the user is required to provide a citable bibcode.
@@ III.2 OTTER API
As discussed in §II.3, some photometry stored in OTTER may be different reductions of the same dataset. This especially affects X-ray and UV data and is intentional because it provides users the freedom to choose their preferred reduction. So, as part of the query implemented in the Python API, we provide an algorithm for finding different reductions of the same data and choosing only one of those photometry points. The algorithm finds any photometry points within a specified date tolerance ($\sim$ 1 day) of the observation (or between the minimum and maximum date for binned data) of the same transient from the same telescope and filter but from a different reference. Then, to choose the photometry to be returned to the user, we first check if only one of the reductions is host subtracted and, if so, we default to that value. If neither reduction or more than one reduction are host subtracted, then we take the most recent reduction. Our testing reveals that this algorithm works very well for the current OTTER dataset. If the user disagrees with this algorithm, they are instead able to “plug in” their favorite algorithm to the query (for example, an algorithm that always prefers the reduction from one paper over another).
@@
IV Current State of the OTTER Dataset
Figure 7: Cumulative histogram of the number of TDE candidates discovered per year, colorized by the wavelength range in which it was discovered. The approximate start date of the major TDE candidate discovery telescopes are labelled on the top. Note the logarithmic scale on the y-axis. Figure 16 is a non-cumulative version of this histogram in Appendix C. A movie of this figure with a linear y-axis is available at: https://github.com/astro-otter/examples/blob/main/tde-discovery-histogram-linear.gif
@@ IV Current State of the OTTER Dataset
Figure 9: Light curves in $\nu L_{\nu}$ space of all of the photometry in OTTER. The colorbars have different minima and maxima based on the number of observations in that wavelength regime and are normalized by the width of the bin in days. Note the two distinct classes of Radio TDE (or TDE candidate) light curves (jetted and non-jetted radio emission). The UV/Optical/IR and X-ray photometry span $\sim 6$ orders of magnitude in luminosity space, demonstrating a diversity of in UVOIR and X-ray TDE candidate luminosities.
@@
Figure 10: Distribution of photometry points in OTTER as a function of Wavelength/Frequency. We have labelled some typical UVOIR and radio filters/bands along the top for reference. The X-ray energy labels are present as a reference and the individual X-ray photometry points are considered solely based on the center wavelength of the energy range used for each observation. Most of the photometry is in the UVOIR bands and the X-ray and Radio distributions have about the same number of photometry points.
@@
As of early December 2025, OTTER contains data on 240 TDE candidates, spanning 34 years of observations, making it the largest catalog of TDE candidate photometry and metadata to date. Figure 7 summarizes the TDE candidate discovery rate by different X-ray, radio, and optical surveys over this time period. It also shows the projection for the TDE discovery rate over the next 5 years as the Rubin Observatory transient discovery rate increases. This is predicted to increase the TDE discovery rate by two orders of magnitude (van Velzen et al., 2011; Bricman & Gomboc, 2020).
@@
For every transient in OTTER, we have at least one set of coordinates. $227/240\approx 95\%$ of the transients in OTTER have a spectroscopic redshift. The transients missing a redshift are MAXI J1807+132, which is an X-ray selected TDE candidate, and the 12 radio selected TDE candidates from Dykaar et al. (2024). Figure 8 shows the redshift distribution of known TDE candidates in OTTER, not corrected for observational bias. The peak appears around $z\sim 0.09$ and 68% of TDE candidates lie in the range from $z\sim 0.03-0.2$, which is likely the result of an observational bias.
@@
$203/240\approx 85\%$ of the transients in OTTER have a discovery date stored. The missing discovery dates are from TDE candidates that were not discovered on a specific date but were rather found in archival searches (e.g., Masterson et al., 2024; Somalwar et al., 2023a; Dykaar et al., 2024). We hesitate to set the discovery date of these events to the first archival detection since, generally, archival searches have much larger gaps between observations than dedicated surveys. This means it is difficult to interpret the date of the first detection as the discovery date under the typical definition.
@@
All the photometry in OTTER is shown in Figures 9, 10, and 11 and contains $\gtrsim$ photometry points. of the TDE candidates in OTTER have at least one photometry point at any wavelength. of the TDE candidates in OTTER have at least one radio observation, $141/240\approx 59\%$ of TDE candidates in OTTER have at least one UVOIR observation, and $82/240\approx 34\%$ of TDE candidates in OTTER have at least one X-ray observation. Some of the missing photometry simply demonstrates inconsistencies in the multiwavelength follow up of TDE candidates. In more recent years, more TDE candidates were followed up more consistently across the electromagnetic spectrum (e.g., Figure 11).
@@
Note that in years since $\sim 2022$, our photometric completeness begins to fall off and we have less photometry after $2024$. This is shown in Figure 11 and is expected because OTTER only contains published photometry, so the more recently discovered TDE candidates simply have fewer published observations.
@@
V Example Applications of OTTER
The OTTER infrastructure has a wide variety of potential applications across the field of transient astronomy, ranging from a quick view of existing public datasets to detailed quantitative comparison with models and locally-stored datasets. Here, we list a few brief examples of applications of the catalog and infrastructure to the TDE candidate population currently stored in OTTER. We provide jupyter notebooks as part of the documentation demonstrating each of these analyses.
V.1 Example 1: Quantifying the Delayed X-ray Emission in Optically-Discovered TDEs
@@
Guolo et al. (2024) found, qualitatively, that the X-ray emission of a subset of optically-selected TDEs is delayed compared to their UV/Optical emission. To further explore this, we quantify the time delay for a large sample of TDEs. We select all of the TDEs in OTTER (as of November 2025) that have , with a few exceptions. We exclude the jetted TDEs Sw J1644+57, Sw J2058+05, and AT 2022cmc since the processes producing the UV/Optical and X-ray emission are unlikely to be the same as the rest of the TDEs in this sample (Zauderer et al., 2011; Levan et al., 2011; Berger et al., 2012; Wiersema et al., 2012; Zauderer et al., 2013; Levan et al., 2016; Yang et al., 2016; Auchettl et al., 2017; Eftekhari et al., 2018; Cendes et al., 2021b; Cenko et al., 2012a; Pasham et al., 2015; Auchettl et al., 2017; Brown et al., 2017). . Note, we do not consider infrared photometry for this analysis ($\lambda_{\rm eff}\geq 1~{\rm\mu m}$) because the infrared emission, which is often dominated by dust echoes, appears to be unrelated to the early-time UV/Optical emission (van Velzen et al., 2021a, 2023). These cuts resulted in a sample that is entirely optically selected.
@@
Since our sample only includes optically discovered TDEs that were later followed up with X-ray observations, there may be a systematic bias of $\Delta t_{\rm del}$ to positive values greater than the true population $\Delta t_{\rm del}$ distribution. We partially account for this bias by introducing upper limits when the maximum was the first observation and lower limits when the maximum was the last observation (e.g., see the arrows in 12a). For subsequent statistical analysis, if $t_{\rm X}$ is an upper limit or $t_{\rm UVO}$ is a lower limit then $\Delta t_{\rm del}$ is treated as a lower limit. In contrast, if $t_{\rm X}$ is a lower limit or $t_{\rm UVO}$ is an upper limit then $\Delta t_{\rm del}$ is treated as an upper limit.
@@
To account for the uncertainties and limits on the delay times in our statistical analysis, we fit a Kaplan-Meier estimator (Kaplan & Meier, 1958) to the $\Delta t_{\rm del}$ distribution [^18] To be more specific, we fit the Kaplan-Meier estimator using the uncertainties and/or limits as interval censored data, essentially treating each of these $\Delta t_{\rm del}$ values as a range (which could have a lower limit of -$\infty$ or an upper limit of $\infty$).. The best-fit Kaplan-Meier estimator has a median $\Delta t_{\rm del}$ of $61$ with 16th and 84th quartiles of 3 and 282 days, respectively. We additionally compute a $95\%$ Greenwood confidence interval (Greenwood, 1926) and find that the true population median of $\Delta t_{\rm del}$ lies in the range 32283-228 days.
Another potential bias, which may also change $\Delta t_{\rm del}$ , occurs because the available X-ray observations for different TDEs are not be equally sensitive. For example, some TDEs have only Swift follow up, but a typical Swift 2ksec observation is insufficient to detect some of the fainter X-ray emission detected with, e.g., XMM-Newton or Chandra (e.g., Guolo et al., 2024). This may result in poor estimates for some of the $t_{\rm X}$ measurements due to a lack of deep X-ray limits. We attempt to account for this by using uncertainties and limits on the measurement of the maximum time, but nevertheless this bias may reduce our constraining power from the data available. Future observations with Einstein Probe (Yuan et al., 2022) will likely be able to reduce this bias and improve the constraining power.
Overall, based on the quartiles of the $\Delta t_{\rm del}$ distribution, we find a conservative upper limit $\Delta t_{\rm del}\lesssim 300$ days, with one exceptional event with an extremely long delay time (PTF09axc [^19] https://otter.idies.jhu.edu/transient/PTF09axc). This indicates that, for most TDEs, we disfavor, but cannot completely rule out, theoretical stream-stream models that predict longer circularization times than a few hundred days. However, as discussed above, there are numerous biases that are difficult to fully account for. Therefore, the data may still be consistent with many of the previously presented models (Piran et al., 2015; Shiokawa et al., 2015; Dai et al., 2018; Metzger, 2022; Steinberg & Stone, 2024).
If the $\Delta t_{\rm del}$ distribution is interpreted as the circularization time, the wide spread we find in $\Delta t_{\rm del}$ could be further evidence that the circularization time is physically related to the initial period of the stellar orbit (Bonnerot & Stone, 2021), SMBH properties (e.g., mass and spin, Steinberg & Stone, 2024), or nodal precession of the plane of the stellar stream (Guillochon & Ramirez-Ruiz, 2013; Bonnerot & Stone, 2020). If, instead, the time delay is interpreted as the dissipation time of a reprocessing envelope, the spread in the distribution can be interpreted as a viewing angle effect (Dai et al., 2018; Metzger, 2022). Moving forward, 1) a more explicit theoretical prediction of this $\Delta t_{\rm del}$ distribution and 2) deeper early-time X-ray observations of TDEs may be necessary to distinguish between them. We defer further investigation of these potential relationships to future work.
V.2 Example 2: The Dusty Environments of TDE Extreme Coronal Line Emitters
@@
As a sample analysis, we plot the peak radio luminosity (a mix of public and private data; Franz, N et al., in prep) against the WISE colors (which we obtain by using the OTTER API to query WISE) for the ECLE-TDEs in Figure 13 and compare these to the non-ECLE-TDEs population. Using the OTTER dataset from November 2025, we find that ECLE-TDEs tend to occur in environments with redder dust than the broader TDE population with a high statistical significance ($p\sim 10^{-7}$, one-sided Kolmogorov-Smirnov test; $p\sim 10^{-5}$, one-sided Wilcoxon rank-sum test). By extension, this may indicate that ECLE-TDEs occur in environments with higher density dust [^21] There is also the possibility that this difference instead indicates a different amount of incident flux onto the dust grains, especially in the case of the archival ECLEs that do not have constraints on the start of the observed emission. But, consider that Hinkle et al. (2024) found that the WISE color is also transient in a subset of our ECLE sample, indicating that the dust heating is transient in these events. We defer further discussion disentangling these scenarios to future work as it will likely require a more detailed multiwavelength investigation.. This is a similar finding to Hinkle et al. (2024) but with a larger sample of events. The radio data in this figure will be discussed in more detail in Franz, N et al., in prep.
V.3 Example 3: Easily Modeling TDE Light-Curves
Another advantage of consolidating the TDE candidate data is that it makes it very easy to apply existing modeling codes, such as MOSFiT (Guillochon et al., 2018; Mockler et al., 2019), redback (Sarin et al., 2024), FitTed (Mummery et al., 2024a), and TDEMass (Ryu et al., 2020a, b). By fitting these models to the data we can derive properties of the SMBH, progenitor star, and cirumnuclear environment. Most notably, these models produce a galaxy-independent measure of the SMBH mass which is either constrained by the light curve as a whole (e.g., MOSFiT) or from the late-time plateau of the light curve (e.g., FitTed).
@@ V.3 Example 3: Easily Modeling TDE Light-Curves
In this work, we focus on applying MOSFiT to some of the objects in the OTTER dataset. The MOSFiT TDE model begins after the circularization ceases, when accretion onto the SMBH is actively occuring. MOSFiT then assumes that the accretion disk produces thermal X-rays which are reprocessed into the UV/Optical. As an example, we can model two TDEs recently discovered by ZTF: TDE 2020zso (Forster et al., 2020a; Ihanec et al., 2020; Hammerstein et al., 2023; Mummery et al., 2024) and TDE 2021sdu (Munoz-Arancibia et al., 2021a; Chu et al., 2021a; Yao et al., 2023; Mummery et al., 2024). We show these models in Figure 14. From MOSFiT we find that the mass of the SMBH (mass of the disrupted star) in TDE 2020zso is $\log_{10}\left(M_{BH}\right)=6.14^{+0.13}_{-0.24}$ ($\log_{10}\left(M_{*}\right)=0.002^{+0.19}_{-0.08}M_{\odot}$) and of TDE 2021sdu is $\log_{10}\left(M_{BH}\right)=5.53^{+0.22}_{-0.25}$ ($\log_{10}\left(M_{*}\right)=0.02^{+0.42}_{-0.18}M_{\odot}$), where the lower and upper uncertainties are the 5% and 95% quantiles, respectively. A jupyter notebook demonstrating how to model these two datasets with MOSFiT is available on the OTTER API documentation webpage and shows that it takes $\lesssim 100$ lines of code to reformat the OTTER data into an input file for MOSFiT.
V.4 Example 4: Validating Optical Transient Classifiers in the Rubin era